OpenStudy (anonymous):

what is the solution to 2cos^2 x+3sinx-3=0?

OpenStudy (jdoe0001):

\(\bf 2{\color{blue}{ cos^2(x)}}+3sin(x)-3=0\implies 2[{\color{blue}{ 1-sin^2(x)}}]+3sin(x)-3=0 \\ \quad \\ 2-2sin^2(x)+3sin(x)-3=0\implies 0=2sin^2(x)-3sin(x)+1\) notice it's just a quadratic, thus solve it likewise

OpenStudy (anonymous):

i mean for the values on the graph between 0 and 360

OpenStudy (mathmale):

You still need to solve the above equation for x. Then, you list all of the x values that lie within the interval [0, 360 degrees). Hint: try solving 2y^2 - 3y + 1 for y. Then apply the same technique to solving the equation in sin x for sin x, first, and for x, second.

OpenStudy (mathmale):

Can you solve 2y^2 - 3y + 1 = 0? Please give this a try.

OpenStudy (mathmale):

You could: 1. factor, or 2. use the Quadratic Formula, or 3. solve for y by graphing

OpenStudy (anonymous):

2y^2 - 3y + 1 = 0. Remind you of 2 Tips in quick solving quadratic equations Tip 1. When a + b + c = 0, one real root is (1), and the other is (c/a) Tip 2. When a - b + c = 0, one real root is (-1), and the other is (-c/a) In the above example (Tip 1), one real root is y1 = 1 and y2 = c/a = 1/2 y1 = cos x = cos 0 and cos 2Pi --> x = .... and ..... y2 = cos x = 1/2 = cos Pi/3 and cos (2Pi - Pi/3) = 5Pi/3-> x = ....., and .....